OpenStudy (anonymous):

plz help Jack's pet dog has been prescribed 40 mg of a medicine per day. Each pill of the medicine has 12 mg of the medicine. Jack has to give his dog a total of 20 pills. If Jack gave his dog medicine for x days, use the expression below to interpret the situation. 12 • 20 − 40x Part A: What does 240 represent? Part B: What does the second term represent? Part C: What does the entire expression represent?

OpenStudy (phi):

Each pill has 12 mg of the medicine total of 20 pills 20 mg per pill times 20 pills is 240 mg (of medicine)

OpenStudy (phi):

can you answer Part A: What does 240 represent?

OpenStudy (anonymous):

240 represents 240 mg. right?

OpenStudy (phi):

yes. but there is one more idea... the total amount of medicine is 240 mg (divided into 20 pills)

OpenStudy (phi):

40 mg of a medicine per day for x days what is 40x (40 mg times x days) ? in other words, what does 40x represent ?

OpenStudy (anonymous):

40mg in x amount of days ?

OpenStudy (phi):

40 mg times 1 day = 40 mg 40 mg times 2 days = 80 mg 40 mg times 3 days = 120 mg what are we keeping track of here ?

OpenStudy (anonymous):

How much medicine is being taken by each day

OpenStudy (phi):

how about, how much medicine we used up ?

OpenStudy (anonymous):

that to also works!

OpenStudy (phi):

What does the entire expression represent? 12 • 20 − 40x the first part 12*20 = 240 is how much we start with 40x is how much we used up (after x days) what are we figuring out here ?

OpenStudy (anonymous):

How many days we would have medicine until we ran out

OpenStudy (phi):

it can't be days. look at the equation 240 mg - 40 mg/day * x days remember that 40x means 40 mg times 1 day = 40 mg 40 mg times 2 days = 80 mg 40 mg times 3 days = 120 mg how much medicine used up. it is so many mg if you start with 240 mg and subtract 40 mg (for example) the answer will be in mg not days. try again. what are we figuring out?

OpenStudy (phi):

you could look at what you get for x=1, 2, 3 240 - 40*1 = 200 240 - 40*2= 160 240 - 40*3= 120 what are we calculating...

OpenStudy (anonymous):

240mg - 40 a day. = 200 We are calculating how much mg we take each day.

OpenStudy (phi):

we know we take 40 mg each day

OpenStudy (anonymous):

how much mg is left?

OpenStudy (phi):

yes, How much medicine in mg we have left after x days For example, after 3 days, we have 120 mg left (we use the formula 240 - 40*3 = 240-120= 120 mg left)

OpenStudy (anonymous):

You the best Phi ! lol

OpenStudy (phi):

yw